how many bricks would it take to skirt a 76x30 foot mobile home if the bricks are 9x2.5 inches
Answer by checkley77(12844)

76*30=2,280 ft^2
9*2.5=22.5 in^2
22.5/12*12=22.5/144=.15625 ft^2 per brick.
2,280/.15625=14,592 bricks will be required.